OpenStudy (saad560):

China's population is 14.46×10814.46×108 . Marshall Islands' population is 602.5×102602.5×102 . How many times larger is China's population than Marshall Islands' population? Enter your answer, using scientific notation, in the boxes. _× 10-

OpenStudy (saad560):

@563blackghost @welshfella @mathmate

OpenStudy (saad560):

@brainzonly

OpenStudy (saad560):

@3mar

OpenStudy (saad560):

@563blackghost

OpenStudy (saad560):

is that so basic

OpenStudy (saad560):

@563blackghost

563blackghost (563blackghost):

Im sorry but im not really good at these kinds of questions sorry maybe @3mar can help :)

OpenStudy (3mar):

Thanks, 563.

OpenStudy (saad560):

@3mar do you know the answer

OpenStudy (3mar):

I will guide you to it, In Sha; Allah

OpenStudy (3mar):

I will guide you to it In Sha' Allah

OpenStudy (3mar):

By the way, did you finish the last question 4⋅(14−3)+124⋅(14−3)+12 ??

OpenStudy (saad560):

because the answer was on mathway

OpenStudy (3mar):

Let me know how would you strat to answer such question?

OpenStudy (3mar):

I will be back with in 15 min. Excise me.

563blackghost (563blackghost):

You should not use Mathway.com since it is basically cheating...unless though you have membership in where it shows on how to solve...

OpenStudy (3mar):

Why is it cheating?

OpenStudy (mathmate):

`14.46×10814.46×108` probably means \(14.46×10814.46×10^8\) Probably that's why no one is able to help you. If you do not write in LaTeX, you can write instead: 14.46×10814.46×10^8 so people know you have ten raised to power 8. Same goes for other numbers.

OpenStudy (3mar):

Thank you, mathmate. You find the trick. @saad560, I think it is solved now. \[\frac{China's population}{ Marshall Islands' population }=\frac{ 14.46*10814.46*10^8 }{ 602.5*102602.5*10^2 }=\frac{ 14.46*10814.46 }{ 602.5*102602.5}*10^6\] \[=2529.64\] SO you can deduce that China's population is about 2530 times larger that Marshall Islands' population